An object projected vertically up from top of a tower took 5 seconds to reach the ground. if the average velocity of the object is 5 metre per second, find its average speed. given gravity is equal to 10 M per second.

Grade:9

2 Answers

ADARSH SINGH
5 years ago
 dear friend ,
 
i think there are a lot of errors in your question and hence , it is impossiblr to solve it.  .. it actually doesnot make any sense to ask average speed by giving average velocity and such an insufficient data . and yes.. to be sure .. accelaration due to gravity can never be equal to 10 M ...PER SECOND.  DO YOU ACTUALLLY BELIEVE THAT ACCELARATION CAN BE 10 MOLAR PER SECOND ?
 
PLEASE REPOST YOUR QUESTION WITH SUFFICIENT DATA AND PROPER LANGUAGE.
 
THANK YOU !
HAVE A NICE DAY!
AND FEEL FREE TO REPLY IN CASE OF ANY QUERIES........
Ashish
5 years ago
  • We have A =10ms^-2
  • ".        " T = 5s
  • average V=5m/s
  • Now, only velocity=50m/s ...(V= u+at)
  • D =s×t
  • 50×5=250m
  • Total distance=250m and total time=5s
  • Average speed =50m/s
  • average speed= 50m/s
  • Thaksss yaar